# Automated Composition of Agents: A Knapsack Approach for Agentic Component Selection

### v1

@@ 5.2 Composing Multi-Agent Systems
For multi-agent benchmarking, we use the end-to-end multi-agent evaluation framework from Shu et al. [30]. Their framework assumes a hierarchical multi-agent collaboration (MAC) architecture where one agent acts as a supervisor and delegates tasks to specialist sub-agents. The benchmarking framework publicly released a MAC benchmarking dataset that includes an inventory of about 20 sub-agents for a few enterprise domains. The evaluation involves using the scenario from the dataset as input and then simulating the conversation between the user and multi-agent team. If the team’s trajectory satisfies the annotated list of assertions, then it is considered a success.
We set up ReAct tool-calling agents as done in the paper [38] and benchmark on two domains: travel and mortgage. We further synthetically augment the original inventory of agents to about 117 agents (Appendix A.6). We arbitrarily set the price of each sub-agent to \$1. All sub-agents use the same underlying model and their tools are being simulated in this setup. In this controlled setting, there is no meaningful variation in runtime or API costs between sub-agents, so assigning a uniform cost allows us to isolate and evaluate the impact of agent selection strategies, making it feasible to compare how this setting differs from Section 5.1 where the pricing was more varied across items. For these experiments, we fix the budget to \$3 and \$6.
For our experiments, we first pass in task descriptions (Appendix A.7), agent inventory, and optional budget to the composer to select the most relevant sub-agents. Then, we set up multi-agent team where the supervisor has the original profile from [30] but with newly selected sub-agents from agent composition. Finally, we simulate the interactions between user and the newly formed MAC team.
@@ Results
Figure 4 shows the results from the multi-agent experiments. We plot overall goal success rate against budget spent on the agent selection. In both domains, online knapsack (\$6) is the highest-performing approach on the pareto frontier. Unlike single-agent experiments, identity no longer shows high success rate as supervisor agent has trouble delegating task to an inventory of more than 100 diverse agents. Possibly, an improved multi-agent framework could be an alternative approach to handle large-scale network of agents [5].
@@ 6 Discussion
Appendix A.9 shows results on several models, including ones from Llama 4 [1] and Qwen 2.5 [32] model families. We observe similar results where online knapsack tends to outperform the baselines. Furthermore, we observe consistency of these results across multiple runs (Table 12) .
@@ A.4 Tool Inventory
We created the tool inventory from Langchain tools and the ToolRet dataset [28]. The tools can be divided into 23 distinct categories based on their functional similarities, such as "personal task assistance", "fact retrieval", and "financial information". The category with the highest number of tools is "personal task assistance", which includes 24 tools. Conversely, categories like "arithmetic calculations," "health information," and "event management" each contain only one tool, indicating a more specialized focus in these areas.

### v2

@@ 5.2 Composing Multi-Agent Systems
For multi-agent benchmarking, we use the end-to-end multi-agent evaluation framework from shu2024towards. Their framework assumes a hierarchical multi-agent collaboration (MAC) architecture where one agent acts as a supervisor and delegates tasks to specialist sub-agents. The benchmarking framework publicly released a MAC benchmarking dataset that includes an inventory of about 20 sub-agents for a few enterprise domains. The evaluation involves using the scenario from the dataset as input and then simulating the conversation between the user and multi-agent team. If the team’s trajectory satisfies the annotated list of assertions, then it is considered a success.
We set up ReAct tool-calling agents as done in the paper [yaoreact] and benchmark on two domains: travel and mortgage. We further synthetically augment the original inventory of agents to about 117 agents (Appendix A.6). We arbitrarily set the price of each sub-agent to \$1. All sub-agents use the same underlying model and their tools are being simulated in this setup. In this controlled setting, there is no meaningful variation in runtime or API costs between sub-agents, so assigning a uniform cost allows us to isolate and evaluate the impact of agent selection strategies, making it feasible to compare how this setting differs from Section 5.1 where the pricing was more varied across items. For these experiments, we fix the budget to \$3 and \$6.
For our experiments, we first pass in task descriptions (Appendix A.7), agent inventory, and optional budget to the composer to select the most relevant sub-agents. Then, we set up multi-agent team where the supervisor has the original profile from [shu2024towards] but with newly selected sub-agents from agent composition. Finally, we simulate the interactions between user and the newly formed MAC team.
@@ Results
Figure 4 shows the results from the multi-agent experiments. We plot overall goal success rate against budget spent on the agent selection. In both domains, online knapsack (\$6) is the highest-performing approach on the pareto frontier. Unlike single-agent experiments, identity no longer shows high success rate as supervisor agent has trouble delegating task to an inventory of more than 100 diverse agents. Possibly, an improved multi-agent framework could be an alternative approach to handle large-scale network of agents [cheninternet].
@@ 6 Discussion
Appendix A.9 shows results on several models, including ones from Llama 4 [meta2024llama4] and Qwen 2.5 [qwen2.5] model families. We observe similar results where online knapsack tends to outperform the baselines. Furthermore, we observe consistency of these results across multiple runs (Table 12) .
@@ A.3 Modifications to ZCL Algorithm for Online Knapsack Composer
We disclose an extended version of the algorithm in Algorithm 2.
Input: Inventory $\mathcal{A}$, task description $x$, budget $B$
Output: Selected components $\mathcal{S}\subseteq\mathcal{A}$
// Initialization
$\mathcal{S}\leftarrow\emptyset$, $\hat{B}\leftarrow B$, $broken\leftarrow\emptyset$ ;
// $\hat{B}$: remaining budget
// Generate skills from task description
$\mathcal{M}\leftarrow\textsc{GenerateSkills}(x)$ ;
// $\{(m_{j},d_{j},Q_{j},w_{j})\}$: name, desc, queries, importance
// Retrieve top-K candidates per skill
foreach skill $m_{j}\in\mathcal{M}$ do
 $\mathcal{T}_{j}\leftarrow\textsc{TopK}(\mathcal{A},d_{j},K)$;
end foreach
// Compute value-to-cost ratio bounds
$L\leftarrow 1/\max_{a\in\mathcal{A}}c_{a}$ ;
// Worst case: max cost, min value
$U\leftarrow\sum_{j}w_{j}/\min_{a\in\mathcal{A}}c_{a}$ ;
// Best case: min cost, max value
// Iterative online knapsack rounds
for $round\leftarrow 1$ to $R$ do
 $covered\leftarrow\emptyset$ ;
 // Reset covered skills each round
 foreach skill $m_{j}\in\mathcal{M}$ do
    foreach component $a_{i}\in\mathcal{T}_{j}$ do
       if $a_{i}\in\mathcal{S}\cup broken$ or $m_{j}\in covered$ or $c_{i}>\hat{B}$ then
          continue ;
          // Skip if processed or infeasible
         end if
       $scores\leftarrow\textsc{TestOnQueries}(a_{i},\mathcal{M})$ ;
       // Test on all skill queries
       if $scores[m_{j}]<0$ then
          $broken\leftarrow broken\cup\{a_{i}\}$ ;
          // Mark as broken
          continue;
         end if
       $v_{i}\leftarrow\sum_{m_{k}\notin covered}w_{k}\cdot\mathbbm{1}[scores[m_{k}]=1]$ ;
       // Value from new skills
       $\rho_{i}\leftarrow v_{i}/c_{i}$ ;
       // Value-to-cost ratio
       $z\leftarrow(B-\hat{B})/B$ ;
       // Fraction of budget spent
       $\Psi\leftarrow\left(\frac{Ue}{L}\right)^{z}\frac{L}{e}$ ;
       // ZCL threshold
       if $\rho_{i}\geq\Psi$ then
          $\mathcal{S}\leftarrow\mathcal{S}\cup\{a_{i}\}$;
          $\hat{B}\leftarrow\hat{B}-c_{i}$;
          $covered\leftarrow covered\cup\{m_{k}:scores[m_{k}]=1\}$;
         end if
      end foreach
   end foreach
end for
return $\mathcal{S}$
Algorithm 2 Online Knapsack Composer - Detailed
@@ A.4 Tool Inventory
We created the tool inventory from Langchain tools and the ToolRet dataset [sculley2015hidden]. The tools can be divided into 23 distinct categories based on their functional similarities, such as "personal task assistance", "fact retrieval", and "financial information". The category with the highest number of tools is "personal task assistance", which includes 24 tools. Conversely, categories like "arithmetic calculations," "health information," and "event management" each contain only one tool, indicating a more specialized focus in these areas.